# Algebraically Skew Embedding of Curves

### v1

@@ 3.3 The class of $\widetilde{\Gamma}$
Let $X\subseteq\mathbb{P}^{N}$ be a degree $d$ genus $g$ smooth curve with its Gauss map $\gamma:X\to{\mathbb{G}\mathrm{r}}(1,N)$ a diffeomorphism to its image. Then
  1. (i)
    for $N=3$,
    $[\widetilde{\Gamma}(X)]=(d^{\vee})^{2}\pi^{*}(\sigma_{21}\otimes\sigma_{21})-j_{*}\left(d^{\vee}\overline{\sigma}_{21}\zeta^{2}+(4d^{\vee}+2g-2)\overline{\sigma}_{22}\zeta\right);$
  2. (ii)
    for $N=4$,
    $[\widetilde{\Gamma}(X)]=(d^{\vee})^{2}\pi^{*}(\sigma_{32}\otimes\sigma_{32})-j_{*}\left(d^{\vee}\overline{\sigma}_{32}\zeta^{4}+(5d^{\vee}+2g-2)\overline{\sigma}_{33}\zeta^{3}\right),$
where $d^{\vee}:=2d+2g-2$.
@@
By Lemma 3.9 we have $[\gamma(X)]=d^{\vee}\sigma_{N-1,N-2}$.
$[E\cap\widetilde{\Gamma}]=$ $d^{\vee}j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)+j_{*}\left((d^{\vee}\overline{\sigma}_{21}(4\overline{\sigma}_{1})+(2g-2)\overline{\sigma}_{22})\zeta^{2}\right)$ (19)
$[E\cap\widetilde{\Gamma}]=$ $(2d+2g-2)j_{*}\left(\overline{\sigma}_{32}\zeta^{5}\right)+j_{*}\left(\left((2d+2g-2)\overline{\sigma}_{32}(5\overline{\sigma}_{1})+(2g-2)\overline{\sigma}_{33}\right)\zeta^{4}\right)$ (20)
The lemma follows from comparing the coefficients above and the coefficients in (12). ∎
3.4 The skew curve in $\mathbb{P}^{3}$
Now that $[\widetilde{D}_{1}],[\widetilde{\Gamma}]$, and the blowup structure are established, we are ready to study the algebraically skew curves in $\mathbb{P}^{3}$.
Theorem 3.11
The twisted cubic is the only algebraically skew curve in $\mathbb{P}^{3}$.
Proof.
Let $X\subseteq\mathbb{P}^{3}$ be a degree $d$ genus $g$ smooth spatial curve, $\widetilde{\Gamma}:=\widetilde{\Gamma}(X)$, and $D_{1}:=D_{1}(2,4)=\mathbb{D}_{1}(1,3)$. We may assume that the curve $X$ is connected and irreducible. Indeed, if a skew spacial curve $X$ has two distinct components $X_{0}$ and $X_{1}$, then their tangent varieties $T_{X_{0}}$ and $T_{X_{1}}$ must intersect as they are both surfaces, which contradicts to the definition of skewness.
We know that a curve $X\subseteq\mathbb{P}^{3}$ is skew if and only if $\Gamma\cap\overline{\mathbb{D}}_{1}=\Delta_{\Gamma}$. As a consequence, we must have $\widetilde{\Gamma}\cap\widetilde{D}_{1}\subseteq\widetilde{\Gamma}\cap E\subseteq E$. By the fact that $\mathsf{dim}\,\widetilde{\Gamma}=2$ and Proposition 2.4, we either have $\widetilde{\Gamma}\cap\widetilde{D}_{1}=\varnothing$ or
$\mathsf{codim}\,(\widetilde{\Gamma}\cap\widetilde{D}_{1})\leq\mathsf{codim}\,\widetilde{\Gamma}+\mathsf{codim}\,\widetilde{D}_{1}=6+1=7.$
Meanwhile, since $\widetilde{\Gamma}\cap E\cong\mathbb{P}(\mathcal{T}_{\gamma(X)})$, we see that $\widetilde{\Gamma}\cap E$ is an irreducible curve as long as the curve $X$ is irreducible. Thus we must have $\widetilde{\Gamma}\cap\widetilde{D}_{1}=\widetilde{\Gamma}\cap E$ if $\widetilde{\Gamma}\cap\widetilde{D}_{1}\neq\varnothing$. This implies $[\widetilde{D}_{1}][\widetilde{\Gamma}]=m[E\cap\widetilde{\Gamma}]$ for some integer $m\geq 0$.
@@ 3.4 The skew curve in $\mathbb{P}^{3}$
$[E\cap\widetilde{\Gamma}]=$ $(2d+2g-2)j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)+(8d+10g-10)j_{*}\left(\overline{\sigma}_{22}\zeta^{2}\right)$ (21)
where $d^{\vee}:=2d+2g-2$. By Lemma 3.5, Lemma 3.10, and Lemma 3.2,
@@
where $O$ is the terms of $\overline{\sigma}_{a}$ with $|a|>1$. Plug this into the equation above, we obtain that
$[\widetilde{D}_{1}][\widetilde{\Gamma}]=$ $d^{\vee}\left(d^{\vee}-2\right)j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)+\left(4{d^{\vee}}^{2}-10d^{\vee}-4g+4\right)j_{*}\left(\overline{\sigma}_{22}\zeta^{2}\right).$ (22)
Notice that $d\geq 3$ as $X$ is nondegenerate, so $d^{\vee}=2d+g-2>0$. Thus, by comparing the coefficients of $j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)$ in (21) and (22), the skewness of $X$ implies $[\widetilde{D}_{1}][\widetilde{\Gamma}]=\left(d^{\vee}-2\right)[E\cap\widetilde{\Gamma}]$. By comparing the coefficients of $j_{*}\left(\overline{\sigma}_{22}\zeta^{2}\right)$ in (21) and (22), we then have
$(d^{\vee}-2)(4d^{\vee}+2g-2)=4{d^{\vee}}^{2}-10d^{\vee}-4g+4,$
which is equivalent to
$02d^{\vee}g=$ 0 .
Since $d^{\vee}>0$, we conclude that $g=0$.
@@
Because the genus zero curve, namely $\mathbb{P}^{1}$, can be parameterized by homogeneous coordinates, it is then an elementary exercise to show that a genus zero spacial skew curve must be the twisted cubic. ∎
@@
Theorem 3.11 can be proved in several different manners, and one may remove assumption of the Gauss map being diffeomorphic in this case. The readers might consider some of these proofs simpler than the one presented above. For instance, one can project the curve to $\mathbb{P}^{1}$, then applies the Hurwitz formula to this projection. We keep the proof above instead of other simpler ones because the presented approach illuminates what to do in higher dimensional cases.
4 The skew curves in $\mathbb{P}^{4}$
Now we move on to the case of a curve $X\subseteq\mathbb{P}^{4}$.
4.1 The excess intersection
@@
To decide whether a given curve $X$ in $\mathbb{P}^{4}$ is skew, we first observe that, due to Proposition 2.4, $[\Gamma]\in\mathsf{A}^{10}(G\times G)$ and $[D_{1}]\in\mathsf{A}^{2}(G\times G)$. Thus, the expected dimension of the intersection $\Gamma\cap D_{1}$ is zero. Whereas we know that the diagonal $\Delta_{\Gamma}$ of $\Gamma:=\gamma(X)\times\gamma(X)$ is a curve lying in the intersection $\Gamma\cap D_{1}$. In addition, in the blowup space $B:=\mathsf{Bl}_{\Delta_{G}}(G\times G)$, the intersection of their proper transforms $\widetilde{\Gamma}\cap\widetilde{D}_{1}$ might not resolve the excess intersection of $\Gamma\cap E$ on the diagonal. In fact, we will see in the following lemma that the intersection $\widetilde{\Gamma}\cap\widetilde{D}_{1}$ contains a double curve supported by $\widetilde{\Gamma}\cap E$.

### v4

@@ 3.4. Self-intersection of one-parameter families of lines in $\mathbb{P}^{3}$
Let $C\subseteq G:={\mathbb{G}\mathrm{r}}(1,3)$ be an irreducible curve. Suppose that $C$ parameterizes a one-parameter family in $\mathbb{P}^{3}$ such that any two distinct lines in the family do not intersect. Then the genus $g$ of the curve $C$ must be 00.
@@
Write $[C]=d^{\vee}\sigma_{21}$, $\Gamma:=C\times C$ and $\widetilde{\Gamma}\subseteq B:=\mathsf{Bl}_{\Delta_{G}}G\times G$ the proper transform of $\Gamma$ as usual. Then two distinct lines in $C$ do not intersect only if, set theoretically, either $\widetilde{\Gamma}\cap\widetilde{D}_{1}=\varnothing$ or $\widetilde{D}_{1}\cap\widetilde{\Gamma}\subseteq E\cap\widetilde{\Gamma}$. Since $C$ is irreducible and $E\cap\widetilde{\Gamma}=\mathbb{P}(\mathcal{T}_{C})$, we must have $[\widetilde{D}_{1}][\widetilde{\Gamma}]=m[E\cap\widetilde{\Gamma}]$ for some integer $m\geq 0$.
@@
(21) $[E\cap\widetilde{\Gamma}]=$ $d^{\vee}j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)+(4d^{\vee}+2g-2)j_{*}\left(\overline{\sigma}_{22}\zeta^{2}\right).$
By Lemma 3.5, Lemma 3.9, and Lemma 3.2,
@@
where $O$ is the terms of $\overline{\sigma}_{a}$ with $|a|>1$. Put this into the equation above, we obtain that
(22) $[\widetilde{D}_{1}][\widetilde{\Gamma}]=$ $d^{\vee}\left(d^{\vee}-2\right)j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)+\left(4{d^{\vee}}^{2}-10d^{\vee}-4g+4\right)j_{*}\left(\overline{\sigma}_{22}\zeta^{2}\right).$
Since by assumption $d^{\vee}\neq 0$, by comparing the coefficients in 21 and 22 we obtain $m=d^{\vee}-2$ and $02d^{\vee}g=$ 0, so we must have $g=0$. ∎
3.5. Self-intersection of scrolls in $\mathbb{P}^{3}$
Before we consider the skew curve in $\mathbb{P}^{3}$, let us examine how we can apply this tool to the classical scroll construction.
Let $C$ be a smooth genus $g$ curve and $\iota_{1},\iota_{2}:C\hookrightarrow\mathbb{P}^{N}$ be two embeddings of $C$ into $\mathbb{P}^{N}$ with degrees $d_{1}$ and $d_{2}$ respectively. Define the scroll $S$ with respect to $(\iota_{1},\iota_{2})$ by
$S:=\bigcup_{t\in C}\left\langle\iota_{1}(t),\iota_{2}(t)\right\rangle.$
In this case, we say that $S$ is a scroll of bidegree $(d_{1},d_{2})$ over $C$. It is convenient to assume a mild generic condition for $S$ a follows. We said that a scroll $S$ with respect to $\iota_{1},\iota_{2}$ is generic if $\iota_{1},\iota_{2}$ are generic under ${\mathbb{P}\mathrm{GL}}(N+1)$ actions. In particular, for a generic $S$ in $\mathbb{P}^{N}$ with $N\geq 3$, the embedded tangent lines $T_{\iota_{1}(t),\iota_{1}(C)}$ and $T_{\iota_{2}(t),\iota_{2}(C)}$ are skew for generic $t\in C$. A scroll $S$ with respect to $\iota_{1},\iota_{2}$ is said to be skew if $\left\langle\iota_{1}(t),\iota_{2}(t)\right\rangle\cap\left\langle\iota_{1}(s),\iota_{2}(s)\right\rangle=\varnothing$ for any two distinct $t,s\in C$. Note that we do not assume that $\iota_{1}(C)$ and $\iota_{2}(C)$ lie in two disjoint linear subspaces in $\mathbb{P}^{N}$ for a scroll.
In this subsection, we answer the question that for what $(d_{1},d_{2})$ and genus $g$ a generic scroll in $\mathbb{P}^{3}$ is skew .
Observe that a scroll $S$ gives us a map $C\to G:={\mathbb{G}\mathrm{r}}(1,3)$ given by $t\mapsto\left\langle\iota_{1}(t),\iota_{2}(t)\right\rangle$. By identifying $C$ with its image in $G$, we have by definition $[C]=(\deg S)\sigma_{21}$. Therefore, to answer the above question it suffices to compute $\deg S$ and then apply Lemma 3.11.
Lemma 3.12
Let $S\subseteq\mathbb{P}^{N}$ be a generic scroll with bidegree $(d_{1},d_{2})$, then $\deg S=d_{1}+d_{2}$.
Proof.
Let $\iota_{1},\iota_{2}$ be defined as above. Recall that $\deg S=d_{1}+d_{2}$ when $\iota_{1}$ and $\iota_{2}$ are embeddings into two disjoint linear subspaces given by two complete linear systems. Indeed, if $\iota_{i}:C\hookrightarrow\mathbb{P}^{N_{i}}\subseteq\mathbb{P}^{N}$ is given by very ample line bundle $\mathcal{L}_{i}$’s with $\mathbb{P}^{N_{1}}\cap\mathbb{P}^{N_{2}}=\varnothing$, then $S$ is the image of the embedding $\mathbb{P}(\mathcal{L}_{1}\oplus\mathcal{L}_{2})\hookrightarrow\mathbb{P}^{N}$ under the hyperplane section bundle, and thus of degree $c_{1}(\mathcal{L}_{1})+c_{1}(\mathcal{L}_{2})=d_{1}+d_{2}$.
The Lemma then follows from the observation that two disjoint curves $C_{1},C_{2}\subseteq\mathbb{P}^{N}$ in general position with fixed degrees can be obtained by projection from $\widetilde{C}_{1},\widetilde{C}_{2}\subseteq\mathbb{P}^{\widetilde{N}}$ with the curves $\widetilde{C}_{1},\widetilde{C}_{2}$ lie in disjoint linear subspaces in $\mathbb{P}^{\widetilde{N}}$. ∎
We now arrive at the following theorem.
Theorem 3.13
Suppose that $S\subseteq\mathbb{P}^{3}$ is a skew generic scroll, then $S$ is a rational normal scroll; that is, $S$ is over $\mathbb{P}^{1}$ with bidegree $(1,1)$.
Proof.
From Lemma 3.11 and Lemma 3.12 we immediately get that $S$ must be over $\mathbb{P}^{1}$.
Treat the embeddings $\iota_{i}:\mathbb{P}^{1}\hookrightarrow\mathbb{P}^{3}$ as $4\times 1$ column vector value functions. The skewness of the scroll $S$ implies that, for $t,s\in\mathbb{P}^{1}$, the matrix $M=(\iota_{1}(t),\iota_{2}(t),\iota_{1}(s),\iota_{2}(s))$ degenerate if and only if $t=s$. Since we have assumed that $S$ is generic, the only solution to $\mathsf{det}\,M=0$ is $t=s$ implies that, up to ${\mathbb{P}\mathrm{GL}}$ actions, $\iota_{i}$’s are of the forms $\iota_{1}(t_{0}:t_{1}):=(t_{0}:t_{1}:0:0)$ and $\iota_{2}(t_{0}:t_{1}):=(0:0:t_{0}:t_{1})$, thus $S$ must be a rational normal scroll. ∎
The following example shows that the generic condition in the above theorem is necessary.
Example 3.14
Let $\iota_{i}:\mathbb{P}^{1}\hookrightarrow\mathbb{P}^{3}$ be defined by
$\iota_{1}(t_{0}:t_{1}):=\begin{pmatrix}t_{0}^{2}\\ 2t_{0}t_{1}\\ t_{1}^{2}\\ 0\end{pmatrix},\qquad\iota_{2}(t_{0}:t_{1}):=\begin{pmatrix}2t_{0}t_{1}\\ t_{1}^{2}\\ 0\\ t_{0}^{2}\end{pmatrix},$
then $\iota_{1},\iota_{2}$ form a skew scroll. The skewness of the scroll can be verified by straightforward computation.
Note that, in the above example, if we set $C_{i}:=\iota_{i}(\mathbb{P}^{1})$, write $P_{i}$ for the plane that contains $C_{i}$ and $L:=P_{1}\cap P_{2}$, then both $C_{1}$ and $C_{2}$ tangent to $L$ at $\iota_{1}(1:0)$ and $\iota_{2}(0:1)$ respectively. The line $\left\langle\iota_{1}(1:0),\iota_{1}(0:1)\right\rangle$ tangent to $C_{1}$ at $\iota_{1}(0:1)$ and $\left\langle\iota_{2}(0:1),\iota_{2}(1:0)\right\rangle$ tangent to $C_{2}$ at $\iota_{2}(1:0)$. Furthermore, the tangent lines $T_{\iota_{1}(t),C_{1}}\cap T_{\iota_{2}(t),C_{2}}\neq\varnothing$ for all $t\in\mathbb{P}^{1}$. In fact, up to ${\mathbb{P}\mathrm{GL}}$ actions, the example above is the only skew scroll of bidegree $(2,2)$. It is unknown to the author if there are skew scrolls of higher bidegrees.
3.6. The skew curve in $\mathbb{P}^{3}$
Lemma 3.11 also helps us to find the algebraically skew curve in $\mathbb{P}^{3}$.
Theorem 3.15
The twisted cubic is the only algebraically skew curve in $\mathbb{P}^{3}$.
Proof.
Let $X\subseteq\mathbb{P}^{3}$ be a degree $d$ genus $g$ smooth spatial curve. We may assume that the curve $X$ is connected and irreducible. Indeed, if a skew spacial curve $X$ has two distinct components $X_{0}$ and $X_{1}$, then their tangent varieties $T_{X_{0}}$ and $T_{X_{1}}$ must intersect as they are both surfaces, which contradicts to the definition of skewness.
@@ 3.6. The skew curve in $\mathbb{P}^{3}$
By applying Lemma 3.11 to the Gauss image of $X$, we see that $X$ being skew implies that $g=0$. Since $\mathbb{P}^{1}$ can be parameterized by homogeneous coordinates, it is then an elementary exercise to show that a genus zero spacial skew curve must be the twisted cubic. ∎
@@
Theorem 3.15 can be proved in several different manners, and one may remove assumption of the Gauss map being diffeomorphic in this case. The readers might consider some of these proofs simpler than the one presented above. For instance, one can project the curve to $\mathbb{P}^{1}$, then applies the Hurwitz formula to this projection. The same result also follows from a more general statement about Terracini loci of curves in $\mathbb{P}^{3}$ [BC23]. We keep the proof above instead of other simpler ones because the presented approach illuminates what to do in higher dimensional cases.
4. The skew curves in $\mathbb{P}^{4}$
Now we move on to one-parameter family of lines and skew curves in $\mathbb{P}^{4}$.
4.1. Self-intersection of one-parameter family of lines in $\mathbb{P}^{4}$
We first consider the general case of a smooth curve $C\subseteq G:={\mathbb{G}\mathrm{r}}(1,4)$. We would like to know under what conditions two distinct lines in the family do not intersect, and if they do intersect, how many pairs of lines intersect.
Set $\Gamma:=C\times C$, then $[\Gamma]\in\mathsf{A}^{10}(G\times G)$. Thanks to Proposition 2.4, we have $[D_{1}]\in\mathsf{A}^{2}(G\times G)$ and the expected dimension of the intersection $\Gamma\cap D_{1}$ is zero. However, we know that the diagonal $\Delta_{\Gamma}$ of $\Gamma$ is a curve lying in the intersection $\Gamma\cap D_{1}$, so we need to resolve the excess intersection on the diagonal.
Following the approach from previous sections, we consider the proper transforms $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ in $B:=\mathsf{Bl}_{\Delta_{G}}G\times G$. This approach works for a generic one-parameter family of lines in the following sense.
Lemma 4.1
Consider a family of lines in $\mathbb{P}^{4}$ parameterized by a smooth curve $C\subseteq G:={\mathbb{G}\mathrm{r}}(1,4)$. Suppose that for a generic point $x\in C$ with a local affine parameterization
$x=\begin{pmatrix}1&0\\ 0&1\\ f&g\end{pmatrix}$
of $C$ around $x$, the column vectors
$\begin{pmatrix}1\\ 0\\ f\end{pmatrix},\begin{pmatrix}0\\ 1\\ g\end{pmatrix},\begin{pmatrix}0\\ 0\\ f^{\prime}\end{pmatrix},\begin{pmatrix}0\\ 0\\ g^{\prime}\end{pmatrix}$
are linearly independent, where $f$ and $g$ are $3\times 1$ column vectors. Then $E\cap\widetilde{D}_{1}\cap\widetilde{\Gamma}$ is not a curve, where $E$ is the exceptional subvariety of the blowup $B:=\mathsf{Bl}_{\Delta_{G}}G\times G$.
4.2. Self-intersection of scrolls in $\mathbb{P}^{4}$
An immediate application of Lemma 4.1 is the self-intersection of generic scrolls in $\mathbb{P}^{4}$.
Theorem 4.2
Let $S$ be a generic bidegree $(d_{1},d_{2})$ scroll over a genus $g$ curve $C$ defined as in previous sections. Set $d^{\vee}:=d_{1}+d_{2}$. Then $S$ is skew if $(d^{\vee})^{2}-(5d^{\vee}+6g-6)=0$. In particular, if a skew generic scroll $S$ with respect to $\iota_{1},\iota_{2}:C\to\mathbb{P}^{4}$ satisfies $T_{\iota_{1}(t),\iota_{1}(C)}\cap T_{\iota_{2}(t),\iota_{2}(C)}=\varnothing$ for all $t\in C$, then $S$ is a rational normal scroll.
Proof.
Let $\iota_{1},\iota_{2}:C\hookrightarrow\mathbb{P}^{4}$ be the embeddings that define the scroll and $\gamma:C\to G:={\mathbb{G}\mathrm{r}}(1,4)$ be given by $x\mapsto\left\langle\iota_{1}(x),\iota_{2}(x)\right\rangle$ as before. Then, Lemma 3.12 gives us $[\gamma(C)]=d^{\vee}\sigma_{32}$, and so, by Lemma 3.9, we have
$[\widetilde{\Gamma}]=(d^{\vee})^{2}\pi^{*}(\sigma_{32}\otimes\sigma_{32})-j_{*}\left(d^{\vee}\overline{\sigma}_{32}\zeta^{4}+(5d^{\vee}+2g-2)\overline{\sigma}_{33}\zeta^{3}\right).$
Meanwhile, from Lemma 3.5 we also have
$[\widetilde{D_{1}}(2,5)]=\pi^{*}(\sigma_{2}\otimes\sigma_{0}+\sigma_{1}\otimes\sigma_{1}+\sigma_{0}\otimes\sigma_{2})-j_{*}(5\overline{\sigma}_{1}+3\zeta).$
Due to the genericity assumption of $S$, we can apply Lemma 4.1 to see that the intersection $\widetilde{D}_{1}\cap\widetilde{\Gamma}$ consists of pairs of intersecting lines in $S$ as well as the points $t\in C$ such that $T_{\iota_{1}(t),\iota_{1}(C)}\cap T_{\iota_{2}(t),\iota_{2}(C)}\neq\varnothing$. Thus the theorem follows from
$\left\llbracket\widetilde{D}_{1}\right\rrbracket\left\llbracket\widetilde{\Gamma}\right\rrbracket=$ $\left\llbracket(d^{\vee})^{2}\pi^{*}(\sigma_{33}\otimes\sigma_{33})-j_{*}\left(5d^{\vee}\overline{\sigma}_{33}\zeta^{5}+3d^{\vee}\overline{\sigma}_{32}\zeta^{6}+3(5d^{\vee}+2g-2)\overline{\sigma}_{33}\zeta^{5}\right)\right\rrbracket$
$=$ $\left((d^{\vee})^{2}-(5d^{\vee}+6g-6)\right)P,$
where $P$ is the class of a point. Indeed, by solving $(d^{\vee})^{2}-(5d^{\vee}+6g-6)=0$ for integers $g\geq 0$ and $d^{\vee}:=d_{1}+d_{2}\geq 2$, we either have $g=0$ with $d^{\vee}=2,3$ or $g=1$ with $d^{\vee}=5$. Since there is no elliptic curves of degree $\leq 2$ in $\mathbb{P}^{4}$, the case of $g=1$ is impossible. Thus we conclude that the only skew generic scrolls in $\mathbb{P}^{4}$ are the rational normal scrolls. ∎
We would like this approach to work for the case of skew curves in $\mathbb{P}^{4}$. However, it turns out that the family of tangent lines of a curve in $\mathbb{P}^{4}$ does not meet the conditions of Lemma 4.1. Thus we need a more careful study of the excess intersection on the diagonal.
4.3. The excess intersection
@@
Given a curve $X\subseteq\mathbb{P}^{4}$. Let $\gamma:X\to G:={\mathbb{G}\mathrm{r}}(1,4)$ be its Gauss map and $\Gamma:=\gamma(X)\times\gamma(X)\subseteq G\times G$. We would like to determine whether $X$ is skew by taking the intersection of the proper transforms $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ in the blowup space $B:=\mathsf{Bl}_{\Delta_{G}}(G\times G)$. The problem is that the intersection $\widetilde{\Gamma}\cap\widetilde{D}_{1}$ might not resolve the excess intersection on the diagonal. In fact, we will see in the following lemma that the intersection $\widetilde{\Gamma}\cap\widetilde{D}_{1}$ contains a double curve supported by $\widetilde{\Gamma}\cap E$.